OpenStudy (anonymous):

What is the inverse of the relation described by y=(2x-3)^2? Please show steps! Will become a fan and give medal to best answer!!

OpenStudy (rosedewittbukater):

Lets see if I remember this.. To find the inverse is to switch the x and y, then solve for y again. I'll make "V" be the square root sign. y = (2x-3)^2 x = (2y-3)^2 2y-3 = Vx 2y = Vx + 3 y = (Vx + 3)/2 The 3 is not part of the square root.

OpenStudy (anonymous):

So only the x on the last three lines are in a square root?

OpenStudy (rosedewittbukater):

Yes

OpenStudy (anonymous):

Thank you!

OpenStudy (rosedewittbukater):

Oh I forgot one thing. Since it's a square root, the answer should be y = (+/- Vx + 3)/2

OpenStudy (anonymous):